# Optimizing Fine-Grained Parallelism Through Dynamic Load Balancing on Multi-Socket Many-Core Systems

### v1

@@ Abstract
Achieving efficient task parallelism on many-core architectures is an important challenge. The widely used GNU OpenMP implementation of the popular OpenMP parallel programming model incurs high overhead for fine-grained, short-running tasks due to time spent on runtime synchronization. In this work, we introduce and analyze three key advances that collectively achieve significant performance gains. First, we introduce XQueue, a lock-less concurrent queue implementation to replace GNU’s priority task queue and remove the global task lock. Second, we develop a scalable, efficient, and hybrid lock-free/lock-less distributed tree barrier to address the high hardware synchronization overhead from GNU’s centralized barrier. Third, we develop two lock-less and NUMA-aware load balancing strategies. We evaluate our implementation using Barcelona OpenMP Task Suite (BOTS) benchmarks. Results from the first and second advances demonstrate up to 1522.8$\times$ performance improvement compared to the original GNU OpenMP. Further improvements from lock-less load balancing show up to 4$\times$ improvement compared to GNU OpenMP using XQueue. Through a rich set of profiling and instrumentation tools, we are able to investigate the runtime behavior of GNU OpenMP and improve its performance on fine-grained tasks by many orders of magnitude.
@@ I Introduction
In the task parallel programming model, computation is broken down into inter-dependent tasks that can be executed concurrently on various cores while respecting data dependencies. Various parallel languages and libraries support this model, such as OpenMP [1], Cilk [2], and Unified Parallel C (UPC) [3]. We focus on OpenMP, a task-centric model in which higher-level parallel constructs such as loops are translated into fine-granularity tasks with dependencies, which the runtime must dynamically schedule to available resources. When a task is enabled by some thread, it is conceptually queued for execution by a future available thread. Unfortunately, OpenMP implementations and their tasking data structures often scale poorly because of excessive use of expensive synchronization operations such as locks to resolve dependencies [4, 5, 6, 7].
Two approaches to avoiding the performance degradation associated with locks are lock-free programming, which typically rely on atomic hardware operations such as compare-and-swap to synchronize without locks, and lock-less programming, which relies on methods for safely manipulating shared data without using locks. We focus here on lock-less programming in which no atomic primitive is used. Prior work has proposed the use of a lock-less, concurrent, multi-producer multi-consumer (MPMC) task queue, called XQueue, in the LLVM-based OpenMP (LOMP) [4]. That work showed significant performance improvements, often delivering 4–6$\times$ speedup, compared to native LLVM OpenMP by reducing the time spent on synchronization operations.
In this paper, we focus on GNU OpenMP (GOMP)—the most common OpenMP implementation that is a part of the mainstream compiler infrastructure GNU Compiler Collection (GCC). We specifically address the unscalable, heavily entangled implementation of GNU OpenMP for fine-grain tasks, overcoming restrictive synchronization such as the excessive use of locks. We then propose, implement, and evaluate lock-less NUMA-aware, dynamic load balancing (DLB) algorithms. The main contributions of this work are as follows:
@@
    We integrate XQueue, a lock-less, relaxed order MPMC task queue, with GOMP, replacing GNU’s unscalable priority task queue and its global task lock. Our benchmark results show that XQueue-enhanced GNU OpenMP (XGOMP) achieves up to 96.5$\times$ performance improvement (for an NQueens application) compared to the original GOMP.
@@
    We propose a new hybrid lock-free (gathering)/lock-less (releasing) distributed tree barrier that yields a theoretical lower bound of half the atomic memory access operations than GNU’s existing barrier which relies on a globally shared atomic task counter. Our approach differs from prior work in LOMP/XLOMP that use LLVM’s default lock-free (not lock-less) barrier.
    We show that this optimization elevates the performance by up to 15.8$\times$ (NQueens) compared to XGOMP, and up to 1522.8$\times$ (NQueens) compared to GOMP.
@@
    To the best of our knowledge, we propose the first lock-less, NUMA-aware dynamic load balancing algorithms for OpenMP tasking. Our approach significantly mitigates load imbalance for both fine-grained and coarse-grained parallelism and achieves significant performance improvements over all previous work. Our approach differs from prior work in XLOMP that relies on a static load balancing (SLB) approach. We also present a comprehensive study of configuration parameters that shows that all benchmarks considered achieve better performance compared to XGOMPTB.
@@
    We design and implement new software profiling tools for GNU to study and optimize how hardware, application, and OpenMP characteristics influence tasking performance that improve on mainstream profiling tools such as Intel VTune and Scalasca. We use these tools to collect statistics such as time spent on task creation, task queue operations, task execution, and per-thread task locality, providing a deeper understanding of performance behavior. We learn that data and task locality plays a critical role in affecting application performance.
@@
The rest of this paper is organized as follows. Section II introduces the motivation and background of this work. Section III describes our integration of XQueue in GNU OpenMP and the design of our distributed tree barrier. Section IV describes our lock-less dynamic load balancing strategies. Section V outlines how we instrument our code to measure performance. Section VI presents a performance evaluation of our work. Section IX reviews related work. Section X summarize our contributions.
@@ II Motivation and Background
The open source GCC is the official compiler for GNU and Linux systems. For OpenMP programs, GCC compiles and links programs with libgomp (GOMP), which implements OpenMP standards but with GNU-specific customizations. LLVM also has an implementation of OpenMP and uses its own compiler, Clang, primarily designed for performance. The performance gap between the GNU and LLVM OpenMP implementations can be significant: see Fig. 1. Here we show execution times for various OpenMP applications from the Barcelona OpenMP Task Suite (BOTS) benchmark [8] when using GNU OpenMP (GOMP), LLVM OpemMP (LOMP), and an implementation of XQueue in LLVM OpenMP (XLOMP). We run these experiments on an Intel Skylake-192 machine, with 192 cores (384 hardware threads). In some benchmarks, GOMP can be $>$1000$\times$ slower than LOMP and $>$4400$\times$ slower than XLOMP. We attribute the GOMP performance issue to its excessive use of synchronization primitives, including locks and atomic operations.
@@ II-A GNU OpenMP (GOMP)
The GOMP runtime library handles OpenMP’s parallel region as a team of worker threads (i.e., workers). An OpenMP team creates and manages a group of OpenMP threads. The team itself is often created under the OpenMP directive parallel. OpenMP wraps platform-specific or user-defined threads, in our experiments we use the pthread [9] library. GNU OpenMP manages tasks, for example, enqueuing and dequeuing tasks, using a single globally shared priority task queue and a child task queue for each task. The global priority task queue maintains the system-wide information, while the child queue maintains child tasks’ dependencies and priorities. GNU OpenMP uses a single global task lock to protect critical regions for task management, scheduling, and other runtime bookkeeping operations. When a worker reaches a scheduling point, it will first acquire the lock before it enters the critical region. When there are no tasks to be executed or scheduled, and current tasks do not have any dependent tasks, the worker exits or waits for other workers to complete.
@@ II-B XQueue
XQueue [4] is a lock-less MPMC, out-of-order queuing mechanism that can scale to 100s of threads with little contention for hardware resources. XQueue uses B-queue, a concurrent Single-Producer-Single-Consumer (SPSC) lock-free queue designed for efficient core-to-core communication. The latency of B-queue operations can be as low as 20 cycles.
@@
Fig. 2 shows XQueue on a 4-core system. Each core has one worker thread with one SPSC queue for each other worker: one Master queue and the others Auxiliary queues. Each item in a queue is a task pointer. Upon task creation, the worker enqueues the task either to its own master queue or to an auxiliary queue of another worker. It applies a round-robin approach across these queues starting with the master queue. If the chosen queue is full, the worker instead executes that task immediately. When dequeuing, the worker first dequeues from the master queue, if no task is found, it then dequeues a task from its auxiliary queues. The dequeued task is also executed immediately.
@@ II-C Load Balancing and Work Stealing
Integrating XQueue with OpenMP is likely to lead to load imbalance as the SLB approach will assign tasks to threads without considering other tasks assigned to those threads. Thus it is important to consider more advanced load balancing techniques to improve performance.
One popular load balancing technique is work stealing. A worker may be either a victim or a thief. A thief is an under-loaded worker who tries to steal work from other overloaded workers. A worker being stolen from is a victim. Typically, pull-based work stealing is used, in which a thief initiates and pulls a task from a victim. However, the queue from which tasks are stolen requires mutual exclusive access as it is accessed by both thief and victim. As a result, traditional work stealing methods do not fit into our lock-less scenario due to their use of synchronization mechanisms like mutexes, spinlocks, and atomic operations to ensure thread safety and correctness [10]. Use of such methods would lead to the runtime wasting more cycles on cache invalidation, cache misses, and excessive main memory access.
A second issue with existing work stealing techniques is that they are not NUMA-aware. We can distinguish between NUMA-local and NUMA-remote data locality on a shared-memory NUMA architecture with respect to a core. Accessing NUMA-local data incurs less latency, and thus cores spend less time waiting for memory operations and higher throughput. Work stealing algorithms must therefore consider whether a thief is to steal from a NUMA-local worker or a NUMA-remote worker, depending on application characteristics.
@@ III-B XGOMPTB: Adding an Efficient Distributed Tree Barrier
We have now removed all excessive atomic operations in GNU OpenMP. The tree barrier combines lock-free and lock-less techniques; it is also easy to understand and implement. To the best of our knowledge, our approach is the first implementation of a hybrid distributed tree barrier for OpenMP tasking with carefully designed lock-less logic. A distributed tree barrier was briefly introduced to LLVM in 2021; however, it was later reverted due to unresolved errors. At this time, there is no distributed barrier in LLVM. LLVM’s barrier employs locks, atomic operations, and does not implement lock-less optimizations.
@@ IV Dynamic Load Balancing
XQueue uses a static round-robin load balancer to distribute tasks among workers. This approach has two limitations: first, load imbalance can occur, and second, data locality in NUMA machines is not considered which may result in slower data access and poor performance.
@@ IV-A Measuring load imbalance
Fig. 3 shows the time spent by each thread in various states for Fibonacci (Fib) and Multisort (Sort) applications. In the left-hand figures, each stacked horizontal bar summarizes one of the 192 threads, with the colors representing, from left to right, threading (purple), tasking (green), task creation (blue), taskwait directive (yellow), barrier (red), and stall (grey).
We focus on tasking (green), task creation (blue), and stall (grey) as threading and barrier are almost indistinguishable in the figure. We consider utilized time as the time spent on tasking (green) and task creation (blue). In the right chart, the horizontal bars represent tasks created (blue) and tasks executed per thread (green). Fib’s Timeline Summary shows that threads with lower thread ID have much shorter utilized time than other threads; in the Task Count Summary, they generated and executed fewer tasks, indicating a clear imbalance in both utilization and task count per thread. We conclude that Fib is poorly balanced in terms of both utilization and task count per thread; For Sort, although the task count appears balanced, mid-range threads have higher utilization, leaving other threads idle and causing performance degradation.
@@ IV-B Lock-less Messaging Protocol
We first need a lock-less messaging protocol for communicating between threads. We extend prior work [10] with a conditionally random victim selection strategy [11], as follows.
We extend the OpenMP-defined thread data structure with two 64-bit memory cells: a round cell and a request cell. The round number is maintained by the victim to track steal requests that have been “handled” enabling potential thieves to check if the victim is able to be stolen from. It is a monotonically increasing number that is initialized to 1 and incremented by the victim each time a steal request is successfully handled. We consider successful handling of a steal request to be when the victim sees the request and processes whether the request is valid. The request cell contains a 40-bit round number and a 24-bit worker ID. The thief sends a request to the victim by writing to the request cell the victim’s round number combined with the thief’s worker ID.
The thief messaging logic is in Alg. 1. In all algorithms, we define $p_{\text{local}}$ as the probability that the thief steals within their local NUMA zone. $tid_{\text{i}}$ is the thread ID of worker $i$, while $ctid_{\text{thief}}$ is that of the current thief. $req$ and $round$ are the current request cell and round number, respectively. The number of tasks to steal per request is $N_{\text{steal}}$. The thief picks a victim randomly, with probability $p_{\text{local}}$ of NUMA-local and $1-p_{\text{local}}$ of NUMA-remote. The thief then compares the victim’s round cell and the round number extracted from the victim’s request cell. If the round number from round cell is greater than the round number from request cell, meaning there has not been a new request for that victim, it then writes a request combining the victim’s current round number from round cell and the thief’s worker ID to the the victim’s request cell, otherwise the request is not sent because there has been a request already.
The thief maintains a timeout counter to deal with long idle times. This counter is incremented each time the thief reaches the same scheduling point while idle and is reset if i) the timeout counter reaches $T_{\text{interval}}$, triggering a retry, or ii) the worker is no longer idle. Timeout can arise if a request is not sent or is overwritten by other thieves, or a victim is idle. When a worker finds a task to execute, it becomes a victim and tries to handle a request if one exists.
@@
We show the victim logic in Alg. 2. A victim first checks its own request. If the round number of its request is equal to its current round number from its round cell, this is considered as a valid request. The victim then processes the request using a DLB strategy (NA-RP or NA-WS) and increments the round number so that it is willing to accept new steal requests.

### v2

@@ Abstract
Achieving efficient task parallelism on many-core architectures is an important challenge. The widely used GNU OpenMP implementation of the popular OpenMP parallel programming model incurs high overhead for fine-grained, short-running tasks due to time spent on runtime synchronization. In this work, we introduce and analyze three key advances that collectively achieve significant performance gains. First, we introduce XQueue, a lock-less concurrent queue implementation to replace GNU’s priority task queue and remove the global task lock. Second, we develop a scalable, efficient, and hybrid lock-free/lock-less distributed tree barrier to address the high hardware synchronization overhead from GNU’s centralized barrier. Third, we develop two lock-less and NUMA-aware load balancing strategies. We evaluate our implementation using Barcelona OpenMP Task Suite (BOTS) benchmarks. We show that the use of XQueue and the distributed tree barrier can improve performance by up to 1522.8$\times$ compared to the original GNU OpenMP. We further show that lock-less load balancing can improve performance by up to 4$\times$ compared to GNU OpenMP using XQueue.
@@ I Introduction
In the task parallel programming model, computation is broken down into inter-dependent tasks that can be executed concurrently on various cores while respecting data dependencies. Various parallel languages and libraries support this model, such as OpenMP [1], Cilk [2], and Unified Parallel C (UPC) [3]. We focus on OpenMP, a task-centric model in which higher-level parallel constructs such as loops are translated into fine-granularity tasks with dependencies, which the runtime must dynamically schedule to available resources. When a task is enabled by some thread, it is conceptually queued for execution by a future available thread. Unfortunately, OpenMP implementations and their tasking data structures often scale poorly due to the excessive use of expensive synchronization operations, such as locks, to resolve dependencies [4, 5, 6, 7].
Two approaches to avoiding the performance degradation associated with locks are lock-free programming, which typically relies on atomic hardware operations such as compare-and-swap to synchronize without locks, and lock-less programming, which relies on methods to safely manipulate shared data without using locks. We focus here on lock-less programming in which no atomic primitives are used. In prior work we proposed the use of a lock-less, concurrent, multi-producer multi-consumer (MPMC) task queue, called XQueue, in the LLVM-based OpenMP (LOMP) [4]. We showed significant performance improvements, often delivering 4–6$\times$ speedup, compared to native LLVM OpenMP by reducing the time spent on synchronization operations.
In this paper, we focus on GNU OpenMP (GOMP)—the most common OpenMP implementation that is a part of the mainstream compiler infrastructure GNU Compiler Collection (GCC). We specifically address the unscalable and entangled implementation of GNU OpenMP for fine-grain tasks, overcoming restrictive synchronization such as the excessive use of locks. We then propose, implement, and evaluate two lock-less NUMA-aware, dynamic load balancing (DLB) algorithms. The main contributions of this work are as follows:
@@
    We integrate XQueue, a lock-less, relaxed order MPMC task queue, with GOMP, replacing GNU’s unscalable priority task queue and global task lock. Our benchmark results show that XQueue-enhanced GNU OpenMP (XGOMP) achieves up to 96.5$\times$ performance improvement (for an NQueens application) compared to the original GOMP.
@@
    We propose a new hybrid lock-free (gathering)/lock-less (releasing) distributed tree barrier that yields a theoretical lower bound of half the atomic memory access operations than GNU’s existing barrier which relies on a globally shared atomic task counter. Our approach differs from prior work in LOMP that uses LLVM’s default lock-free (not lock-less) barrier. We show that this optimization (XGOMPTB) improves the performance by up to 15.8$\times$ (NQueens) compared to XGOMP, and up to 1522.8$\times$ (NQueens) compared to GOMP.
@@
    To the best of our knowledge, we propose the first lock-less, NUMA-aware dynamic load balancing algorithms for OpenMP tasking. Our approach significantly mitigates load imbalance for both fine-grained and coarse-grained parallelism and achieves significant performance improvements over all previous work. We also present a comprehensive study of configuration parameters that shows that all benchmarks considered achieve better performance compared to XGOMPTB.
@@
    We design and implement new software profiling tools for GNU to study and optimize how hardware, application, and OpenMP characteristics influence tasking performance. We use these tools to collect statistics such as time spent on task creation, task queue operations, task execution, and per-thread task locality, providing a deeper understanding of performance behavior. We learn that data and task locality plays a critical role in affecting application performance.
@@
The rest of this paper is organized as follows. Section II introduces the motivation and background of this work. Section III describes our integration of XQueue in GNU OpenMP and the design of our distributed tree barrier. Section IV describes our lock-less dynamic load balancing strategies. Section V outlines how we instrument our code to measure performance. Section VI presents a performance evaluation of our work. Section VII presents a real-world application that benefits from our work. Section VIII discusses how people should tune our DLB for optimal performance. Section IX reviews related work. Section X summarizes our contributions.
@@ II Motivation and Background
The open source GCC is the official compiler for GNU and the standard compiler for many Linux distributions. For OpenMP programs, GCC compiles and links programs with libgomp (GOMP), which implements OpenMP standards but with GNU-specific customizations. LLVM also has an implementation of OpenMP and uses its own compiler, Clang, primarily designed for performance. The performance gap between the GNU and LLVM OpenMP implementations can be significant: see Fig. 1. Here we show execution times for various OpenMP applications from the Barcelona OpenMP Task Suite (BOTS) benchmark [8] when using GNU OpenMP (GOMP), LLVM OpemMP (LOMP), and an implementation of XQueue in LLVM OpenMP (XLOMP). We run these experiments on an Intel Skylake-192 machine, with 192 cores (384 hardware threads). In some benchmarks, GOMP can be $>$1000$\times$ slower than LOMP and $>$4400$\times$ slower than XLOMP. We attribute the GOMP performance issue to its excessive use of synchronization primitives, including locks and atomic operations.
@@ II-A GNU OpenMP (GOMP)
The GOMP runtime library handles OpenMP’s parallel region as a team of worker threads (i.e., workers). An OpenMP team creates and manages a group of OpenMP threads. The team itself is often created under the OpenMP directive parallel. OpenMP wraps platform-specific or user-defined threads. In our experiments, we use the pthread [9] library. GNU OpenMP manages tasks, for example, enqueuing and dequeuing tasks, using a single globally shared priority task queue and a child task queue for each task. The global priority task queue maintains the system-wide information, while the child queue maintains child tasks’ dependencies and priorities. GNU OpenMP uses a single global task lock to protect critical regions for task management, scheduling, and other runtime bookkeeping operations. When a worker reaches a scheduling point, it will first acquire the lock before it enters the critical region. When there are no tasks to be executed or scheduled, and current tasks do not have any dependent tasks, the worker exits or waits for other workers to complete.
@@ II-B XQueue
XQueue [4] is a lock-less MPMC, out-of-order queuing mechanism that can scale to hundreds of threads with little contention for hardware resources. XQueue uses B-queue, a concurrent Single-Producer-Single-Consumer (SPSC) lock-free queue designed for efficient core-to-core communication. The latency of B-queue operations can be as low as 20 cycles.
@@
Fig. 2 illustrates how XQueue is used on a 4-core system. Each core has one worker thread with a single SPSC Master queue and one Auxiliary SPSC queue for each other worker. Each item in a queue is a task pointer. Upon task creation, the worker enqueues the task to either its own master queue or to an auxiliary queue of another worker. It applies a round-robin approach across these queues starting with the master queue. If the chosen queue is full, the worker instead executes that task immediately. When dequeuing, the worker first dequeues from the master queue, if no task is found, it then dequeues a task from its auxiliary queues. The dequeued task is also executed immediately.
@@ II-C Load Balancing and Work Stealing
Integrating XQueue with OpenMP is likely to lead to load imbalance as the static load balancing (SLB) approach will assign tasks to threads without considering other tasks assigned to those threads. Thus, it is important to consider more advanced load balancing techniques to improve performance.
One popular load balancing technique is work stealing. A worker may be either a victim or a thief. A thief is an underloaded worker who tries to steal work from other overloaded workers. A worker being stolen from is a victim. Typically, pull-based work stealing is used, in which a thief initiates and pulls a task from a victim. However, the queue from which tasks are stolen requires mutual exclusive access as it is accessed by both thief and victim. As a result, traditional work stealing methods do not match our lock-less scenario due to their use of synchronization mechanisms, such as mutexes, spinlocks, and atomic operations, to ensure thread safety and correctness [10]. Use of such methods would lead to the runtime wasting more cycles on cache invalidation, cache misses, and excessive main memory access.
A second issue with existing work stealing techniques is that they are not NUMA-aware. In a NUMA architecture, we can distinguish between data being either NUMA-local and NUMA-remote with respect to a core. Accessing NUMA-local data incurs less latency, and thus cores spend less time waiting for memory operations and can achieve higher throughput. Work stealing algorithms must therefore consider whether a thief is to steal from a NUMA-local worker or a NUMA-remote worker, depending on application characteristics.
@@ III-B XGOMPTB: Adding an Efficient Distributed Tree Barrier
With these modifications, we have now removed all excessive atomic operations in GNU OpenMP. The tree barrier combines lock-free and lock-less techniques; it is also easy to understand and implement. To the best of our knowledge, our approach is the first implementation of a hybrid distributed tree barrier for OpenMP tasking with carefully designed lock-less logic. A distributed tree barrier was briefly introduced to LLVM in 2021; however, it was later reverted due to unresolved errors. LLVM’s barrier implementation uses locks, atomic operations, and does not implement lock-less optimizations.
@@ IV Dynamic Load Balancing
XQueue uses a static round-robin load balancer to distribute tasks among workers. This approach has two limitations: first, load imbalance can occur, and second, data locality in NUMA machines is not considered, which may result in slower data access and poor performance.
@@ IV-A Measuring load imbalance
Fig. 3 shows the time spent by each thread in various states for Fibonacci (Fib) and Multisort (Sort) applications. In the left-hand figures, each stacked horizontal bar summarizes one of the 192 threads, with the colors representing, from left to right, tasking (green), task creation (blue), taskwait directive (yellow), barrier (red), and stall (grey).
We focus on tasking (green), task creation (blue), taskwait directive (yellow) and stall (grey) as barrier (red) is almost indistinguishable in the figure. We consider utilized time as the time spent on tasking (green) and task creation (blue). In the right chart, the horizontal bars represent tasks created (blue) and tasks executed per thread (green). Fib’s Timeline Summary shows that threads with lower thread ID have much shorter utilized time than other threads; in the Task Count Summary, they generated and executed fewer tasks, indicating a clear imbalance in both utilization and task count per thread. We conclude that Fib is poorly balanced in terms of both utilization and task count per thread. For Sort, although the task count appears balanced, mid-range threads have higher utilization, leaving other threads idle and causing performance degradation.
@@ IV-B Lock-less Messaging Protocol
Implementing lock-less DLB strategies first requires a lock-less messaging protocol for communicating between threads. We extend our prior work [10] with a conditionally random victim selection strategy [11], as follows.
We extend the OpenMP-defined thread data structure with two 64-bit memory cells: a round cell and a request cell. The round number is maintained by the victim to track steal requests that have been “handled” enabling potential thieves to check if the victim is able to be stolen from. It is a monotonically increasing number that is initialized to 1 and incremented by the victim each time a steal request is successfully handled. We consider successful handling of a steal request to be when the victim sees the request and processes whether the request is valid. The request cell contains a 40-bit round number and a 24-bit worker ID. The thief sends a request to the victim by writing the victim’s round number combined with the thief’s worker ID to the request cell.
The thief messaging logic is in Alg. 1. In all algorithms, we define $p_{\text{local}}$ as the probability that the thief steals within their local NUMA zone. $tid_{\text{i}}$ is the thread ID of worker $i$, while $ctid_{\text{thief}}$ is that of the current thief. $req$ and $round$ are the current request cell and round number, respectively. The number of tasks to steal per request is $N_{\text{steal}}$. The thief picks a victim randomly, with probability $p_{\text{local}}$ of NUMA-local and $1-p_{\text{local}}$ of NUMA-remote. The thief then compares the victim’s round cell and the round number extracted from the victim’s request cell. If the round number from round cell is greater than the round number from request cell, meaning there has not been a new request for that victim, it then writes a request combining the victim’s current round number from round cell and the thief’s worker ID to the the victim’s request cell; otherwise the request is not sent because there has been a request already.
The thief maintains a timeout counter to deal with long idle times. This counter is incremented each time the thief reaches the same scheduling point while idle and is reset if i) the timeout counter reaches $T_{\text{interval}}$, triggering a retry, or ii) the worker is no longer idle. Timeout can arise if a request is not sent or is overwritten by other thieves, or if a victim is idle. When a worker finds a task to execute, it becomes a victim and tries to handle a request if one exists.
@@
We show the victim logic in Alg. 2. A victim first checks its own request. If the round number of its request is equal to its current round number from its round cell, this is considered a valid request. The victim then processes the request using a DLB strategy (NA-RP or NA-WS) and increments the round number so that it is willing to accept new steal requests.